# Charting the main sequence of star-forming galaxies out to redshifts $z\sim 7$

### v1

@@ 5.2 Temperature-redshift relation
The evolution of the dust temperature with redshift has been debated numerous times in the literature. There are three main sets of claims that can be distinguished – those that argue that, at fixed $L_{\rm IR}$, $T_{\rm d}$ is lower at higher redshifts (e.g. Chapman et al. 2002; Pope et al. 2006; Magnelli et al. 2014), those that claim $T_{\rm d}$ to be increasing with $z$ (e.g. Schreiber et al. 2018; Faisst et al. 2020), and those that show no evolution at all (e.g. Dudzevičiūtė et al. 2020; Reuter et al. 2020; Drew & Casey 2022). As shown by Drew & Casey (2022), the apparent increase of the dust temperature at lower redshift galaxies is often caused by the bias, whereby colder systems had been fit with warmer-temperature SEDs due to the lack of long wavelength data (and hence no photometric constraints on the Rayleigh-Jeans tail of the dust emission curve). A number of works that find no evolution of $T_{\rm d}$ with redshift are based on brightest sources selected at (sub-)mm wavelengths, not necessarily representative of the bulk of the star-forming galaxy population on the main sequence (e.g. Dudzevičiūtė et al. 2020; Reuter et al. 2020). Drew & Casey (2022) find that for the sample of $z<2$ galaxies the IR luminosity anti-correlates with the dust emission rest-frame peak wavelength (and hence correlates with the luminosity-weighted dust temperature). They also determine that the $z<2$ $T_{\rm d}$-$L_{\rm IR}$ relation does not evolve with redshift. Given that the star formation rates (and $L_{\rm IR}$) for star-forming main-sequence galaxies, at fixed stellar mass, increase with redshift, the redshift-independent correlation of $L_{\rm IR}$ and $T_{\rm d}$ produces hotter SEDs at higher $z$’s.
Figure 5: The best-fit star-forming main sequence shown at each redshift bin studied in this work as the black solid line. For comparison, we show the relations of Speagle et al. (2014), Tomczak et al. (2016) and Popesso et al. (2023), plotted with blue dashed, green dash-dotted and red dotted lines, respectively. While all the functions tend to roughly agree with each other, some inconsistencies can clearly be seen, discussion of which is presented in Section 5.3.
@@
In the left panel of Figure 4 we plot the peak dust temperatures (derived from best-fit SEDs shown in Figure 11 using Wien’s law) for the star-forming sample of galaxies as a function of the IR luminosity, color-coded with redshift. We can see that, as in the case of Drew & Casey (2022), $T_{\rm d}$ clearly correlates with $L_{\rm IR}$ (even when confining our sample to sources at $z<2$). However, at any given redshift, we do not find any evolution of the dust temperature with $L_{\rm IR}$. If anything, similarly to Schreiber et al. (2018), we can see a minor decline of $T_{\rm d}$ at highest IR luminosities at $z\lesssim 2$. As discussed in Liang et al. (2019), the differences in dust temperatures are caused mainly by the variations of the specific star formation rate (sSFR; see the right panel of Figure 6). In this scenario, higher sSFR systems have more young ($t\lesssim 10$ Myrs) star clusters, which heat the dust to higher temperatures, producing more emission in mid-IR wavelengths and hence increasing the values of the effective peak temperatures. This is consistent with the slight decrease of $T_{\rm d}$ at highest stellar masses and the increase with redshift seen in the left panel of Figure 4. Since, at any given redshift, the dust temperature variations are small, we plot the mean values of $T_{\rm d}$ as a function of $z$ (black dots) in the right panel of Figure 4. A clear linear relation can be seen, with the best-fit empirical formula:
$T^{\rm MS}_{\rm d}/{\rm K}=(6.0\pm 0.3)\times z+(17.8\pm 1.0)$ (9)
@@
shown as a black solid line. For comparison, the results of Schreiber et al. (2018), Viero et al. (2022) and Ismail et al. (2023) are shown as red dotted, green dash-dotted and blue dashed lines, respectively. It can be seen that our results, as well as the relation of Schreiber et al. (2018), give slightly higher temperatures at low redshifts than Ismail et al. (2023). This is most likely caused by the fact that both in this work and in the work of Schreiber et al. (2018), the emissivity index, $\beta$, has been fixed at constant value (1.96 and 1.5, respectively). As shown in Ismail et al. (2023), $\beta$ correlates with dust temperature, where warmer SEDs are characterized by lower emissivity values. At low redshifts, where $T_{\rm d}$ is $\sim 20$ K, $\beta$ is expected to vary between 2.5 and 3.5. Fixing $\beta$ at lower values will naturally increase the best-fit dust temperatures. The effect is stronger for lower assumed values of emissivity, which can be seen in the figure. It is also most obvious for low redshifts, as at higher $z$’s, where the dust temperature increases, the emissivity approaches the values assumed in this work. The $T_{\rm d}$-$z$ relation determined by Viero et al. (2022) increases significantly faster than the remaining curves with redshift. This is simply caused by the different definition of the dust temperature in Viero et al. (2022), where Casey (2012) dust emission SEDs were used with effective temperature $T^{\rm eff}_{\rm d}$. Different temperature definition is necessary because, due to the interstellar dust in galaxies exhibiting a range of temperatures, the Casey (2012) dust emission curve has a power law component at mid-IR wavelengths. As explained in Section 4.1, the power-law component at mid-IR gives better fits to Herschel PACS photometry. Translating from $T^{\rm eff}_{\rm d}$ to $T_{\rm d}$ for the relationship found by Viero et al. (2022) produces a curve which is consistent with the findings of Ismail et al. (2023).
5.3 Comparison to Literature
@@
In Figure 5 we show the comparison of the SFR-$M_{\ast}$ relation for star-forming galaxies (star-forming main sequence) found in this work (black solid lines) with those found by Speagle et al. (2014), Tomczak et al. (2016) and Popesso et al. (2023), plotted with blue dashed, green dash-dotted and red dotted lines, respectively. The lines represent the best-fit functions at the central value of each redshift bin studied in this work. The first feature that can be seen is the lack of the turnover for the SFR-$M_{\ast}$ relation of Speagle et al. (2014). As pointed out by Popesso et al. (2023), this was likely caused by the fact that the sample studied in Speagle et al. (2014) was mostly confined to sources with stellar masses below $10^{11}\,{\rm M_{\odot}}$.
@@
At $z\gtrsim 2$ and ${\rm log}(M_{\ast}/{\rm M_{\odot}})\gtrsim 11$, our values of SFR tend to be larger than those of Popesso et al. (2023). The main reason for this is that the determination of the high-mass end of the main sequence at high redshifts is, in both cases, based on a handful of sources. Looking at Figure 3 of Popesso et al. (2023), it can be seen that at the highest redshift both the bending and the normalization of the main sequence are underpredicted by the adopted functional forms. The errors on the main sequence parameters at high redshifts are significant enough for our best-fit function to be consistent with the individual data points of Popesso et al. (2023). In addition, in most of the works compiled by Popesso et al. (2023), the IR luminosities, and hence SFRs, were determined from near- and mid-IR data fit with relatively cold SEDs ($\sim 30$ K as opposed to $\sim 50$ K peak temperatures found at high-$z$’s in this work – Figure 4), which can bias the results towards lower values of SFR. On the other hand, the presence of starburst galaxies can overpredict the high-mass end of the main sequence at high redshifts. At low redshifts, for ${\rm log}(M_{\ast}/{\rm M_{\odot}})\lesssim 10.5$, our values of the SFR are lower than those of Tomczak et al. (2016) and Popesso et al. (2023). This is, again, most likely caused by fitting the IR data with warm SEDs, where dust temperatures higher than $\sim 20$ K were assumed.
@@
Figure 6: The logarithm of the specific star formation rate, derived from the best-fit functional forms of the star-forming main sequence found in this work, plotted as a function of redshift and color-coded with stellar mass. The shape of the relations is consistent with the redshift evolution of the molecular gas fraction in star-forming galaxies, as discussed in Section 5.3. For comparison we show the findings of Popesso et al. (2023) at two stellar mass regimes. The inset plot shown the relation between the sSFR and peak dust temperature. The similar shapes of the sSFR evolution with $z$ and $T_{\rm d}$ is consistent with the scenario, presented in Liang et al. (2019), in which different dust temperature are mainly driven by the variations in the specific star formation rate (see Section 5.2).
@@
In Figure 6 we show the corresponding evolution of the specific star formation rate with redshift, color-coded with the stellar mass. A clear turnover can be seen at $z\sim 2$, which is a direct consequence of the bending of the normalization of the main sequence with redshift (Equation 8). This, on the other hand, as explained in Section 5.1, is likely caused by the increase in molecular gas fraction, found by the Wang et al. (2022) to be exponentially increasing with redshift. The sSFR-$z$ relation exhibits lower normalization at higher stellar masses, since at masses $\gtrsim M_{0}$ the main sequence flattens, decreasing the resulting sSFR. In the inset plot the evolution of the sSFR with the peak dust temperature can be seen. Similar shapes of the sSFR relation with both redshift and dust temperature are consistent with the suggestion (Liang et al. 2019), that the increase of the dust temperature with redshift, found in Section 5.2, is indeed caused by the corresponding increase in the sSFRs.
@@
For comparison, relation of Popesso et al. (2023) for ${\rm log}(M_{\ast}/{\rm M_{\odot}})\simeq 9.5$ and 11.0 are shown as black dotted and dashed lines, respectively. The inconsistencies at low and high redshifts come from different normalizations of the corresponding main sequences, as explained above. In order to investigate those further, in the following section we explore how the evolution of the main sequence found in this work and the works of Speagle et al. (2014), Tomczak et al. (2016) and Popesso et al. (2023) affects the ‘growth’ of the galaxy stellar mass function (GSMF).
@@ 5.4 Growth of the Galaxy Stellar Mass Function
Figure 7: The comparison between the observed and simulated star-forming galaxy stellar mass functions (GSMFs). The simulated GSMFs are constructed by ‘growing’ the observed function of McLeod et al. (2021), between the redshifts listed in the upper right corner of each panel, according to Equation 10. The mass evolution resulting from assumed star-forming main sequences of Speagle et al. (2014), Tomczak et al. (2016) and Popesso et al. (2023), are plotted with blue dashed, green dash-dotted and red dotted lines, respectively. The simulated function constructed using the results of this work are presented in solid black, while the thick solid grey line represents the observed GSMF at the final redshift of each simulation. The inconsistencies at low- and high-mass ends of the simulated GSMFs come from differences in the corresponding shapes of the MSs involved (Figure 5). The significant shift in densities between the observed and simulated GSMFs is caused by the fraction of star-forming galaxies, simulated between the redshifts in question, entering a quiescent mode, as discussed in Section 5.4.
@@
Figure 8: The comparison between the observed and simulated galaxy stellar mass functions (GSMFs), where both star-forming and quiescent galaxies are considered. To extend the redshift range, the simulated GSMFs are constructed by ‘growing’ the observed function of Davidzon et al. (2017), between the redshifts listed in the upper right corner of each panel, according to Equation 10. The mass evolution results from assuming the SFR-$M_{\ast}$ relation characteristic of galaxy samples including both the star-forming and quiescent sources, where we compare our results, shown in solid black, with those of Tomczak et al. (2016), depicted with green dash-dotted line. The thick solid grey line represents the observed GSMF at the final redshift of each simulation. It can be seen that the significant gap in densities, apparent in Figure 7, is no longer present, indicating that, indeed, the inconsistencies were caused by the fraction of sources evolving into quiescent mode, as discussed in Section 5.4. The low-mass slope of simulated GSMFs at $z\gtrsim 2$ is inconsistent with the observed function, indicating that the corresponding slope of the SFR-$M_{\ast}$ relations most likely evolves towards steeper values towards earlier epochs.
@@
The star-forming main sequence effectively quantifies the stellar mass growth of a galaxy with time. We can, therefore, compare the evolution of the galaxy stellar mass function as implied by any given main sequence with the observed one. The procedure is described in detail in Leja et al. (2015) and can be summarized as follows. At each redshift, the stellar mass for a given GSMF is being ‘grown’ assuming the SFR from the main sequence, causing the GSMF to shift towards higher masses. Additionally, mainly due to winds and outflows, a fraction of the mass is ejected from a stellar population, a careful treatment of which has been presented in Moster et al. (2013). The stellar mass growth with time can, therefore, be expressed as:
@@
We start the procedure with the observed star-forming stellar mass function of McLeod et al. (2021). The stellar mass at any given redshift is evolved according to Equation 10 assuming different star-forming main sequences. In Figure 7 we show the results of this exercise, with the observed star-forming GSMF at the final redshift shown in grey. The stellar mass function, produced by evolving the observed GSMF between redshifts listed in the upper right corner of each panel, for the MS of Speagle et al. (2014), Tomczak et al. (2016) and Popesso et al. (2023) are plotted with blue dashed, green dash-dotted and red dotted lines, respectively. The results from this work are shown with black solid line. It can be seen that the evolved GSMFs differ at both low and high-mass ends, which is due to the differences in the shapes of the corresponding main sequences. At low redshifts and stellar masses the MS found in this work predicts lower values of the SFR, which causes the GSMF to ‘grow’ more slowly in mass and hence the shift towards higher masses is in this case smaller than those produced by Speagle et al. (2014), Tomczak et al. (2016) and Popesso et al. (2023). Similarly, at high redshift and stellar masses, the stellar mass shifts are slightly larger (except for the case of Speagle et al. 2014, where no bending was assumed). The most striking feature of Figure 7 is the inconsistency of all the evolved functions with the observed GSMF. There are two main reasons for that. Firstly, the mass evolution of the GSMF is going to be affected by galaxy mergers, where mergers are expected to decrease the number density and contribute to the mass growth of galaxies. As shown by Leja et al. (2015), however, including mergers will change the evolved GSMF at low masses by only $0.1-0.2$ dex between redshifts 2.25 and 0.5. The second reason is the fact that, as GSMF evolves in time, a fraction of star-forming galaxies will transform into passive sources, effectively shifting the evolved stellar mass function down towards lower densities.
@@
In order to account for this effect, we additionally evolve the GSMF, where both star-forming and passive galaxies have been included. In order to extend this exercise to higher redshifts, we adopt the stellar mass function of Davidzon et al. (2017), where $z\simeq 6$ has been reached. The SFR-$M_{\ast}$ relation used in this case includes all galaxies (star-forming and quiescent) and is shown in the left panel of Figure 3. The resulting evolved GSMF at different redshifts is compared to the observed one in Figure 8 (black thin and grey thick lines, respectively). Since Speagle et al. (2014) and Popesso et al. (2023) did not publish the SFR-$M_{\ast}$ relations for the galaxy samples that include the quiescent sources, we compare our results with Tomczak et al. (2016) only (green dashed line). It is clear from Figure 8 that the obvious shift in galaxies number density between the observed and simulated functions, apparent in Figure 7, is no longer present. This confirms that the inconsistency in the star-forming GSMFs is caused by the decrease in density due to the fraction of star-forming galaxies transforming into passive sources. For low masses at $z\gtrsim 2$ the low-mass end of the simulated stellar mass functions tends to be steeper than the observed ones. This is caused by the low-mass slope of the observed GSMF undergoing the evolution in redshift towards higher values. In order for our SFR-$M_{\ast}$ relation to produce results consistent with this, the low-mass slope, $\gamma$, from Equation 6 would also need to evolve. However, as explained in Section 4.2, since our data does not probe low stellar masses at high redshifts, $\gamma$ was assumed to be constant. The fact that the constant $\gamma$ gives steeper low-mass slopes of the GSMF at high $z$’s indicates that the power-law slope of the SFR-$M_{\ast}$ relation for all galaxies (star-forming and quiescent) should increase in redshift. Similarly, since the low-mass slope of the star-forming GSMF, found in both Davidzon et al. (2017) and McLeod et al. (2021), increases at $z\gtrsim 2$, the slope of the star-forming main sequence should also evolve. At $z\gtrsim 3$ the high-mass end of the simulated GSMF overpredicts the number of galaxies. While the errors on the observed high-redshift stellar mass function at these masses are significantly large ($\sim 1$ dex), it is also possible that the high-mass end of the SFR-$M_{\ast}$ relations in this work, and hence the growth of the stellar masses, is contaminated by the presence of starburst galaxies. Finally, it can be seen in Figure 8 that at $z\lesssim 2$ the SFR-$M_{\ast}$ all-galaxies relation of Tomczak et al. (2016) slightly overpredicts the resulting GSMFs at low masses, most likely indicating that the normalization of the main sequence at low $z$ is in this case too high.

### v2

@@ 5.2 Temperature-redshift relation
The evolution of the dust temperature with redshift has been debated numerous times in the literature. There are three main sets of claims that can be distinguished – those that argue that, at fixed $L_{\rm IR}$, $T_{\rm d}$ is lower at higher redshifts (e.g., Chapman et al. 2002; Pope et al. 2006; Magnelli et al. 2014), those that claim $T_{\rm d}$ to be increasing with $z$ (e.g., Schreiber et al. 2018; Faisst et al. 2020), and those that show no evolution at all (e.g., Dudzevičiūtė et al. 2020; Reuter et al. 2020; Drew & Casey 2022). As has been shown by Drew & Casey (2022), the apparent increase in the dust temperature in lower-redshift galaxies is often caused by the bias, whereby colder systems had been fit with warmer-temperature SEDs due to the lack of long wavelength data (and hence no photometric constraints on the Rayleigh-Jeans tail of the dust emission curve). A number of works that have found no evolution of $T_{\rm d}$ with redshift are based on brightest sources selected at (sub-)mm wavelengths, not necessarily representative of the bulk of the star-forming galaxy population on the MS (e.g., Dudzevičiūtė et al. 2020; Reuter et al. 2020). Drew & Casey (2022) have found that for the sample of $z<2$ galaxies, the IR luminosity anti-correlates with the dust emission rest-frame peak wavelength (and hence correlates with the luminosity-weighted dust temperature). They have also determined that the $z<2$ $T_{\rm d}$-$L_{\rm IR}$ relation does not evolve with redshift. Given that the SFRs (and $L_{\rm IR}$) for star-forming MS galaxies, at fixed stellar mass, increase with redshift, the redshift-independent correlation of $L_{\rm IR}$ and $T_{\rm d}$ produces hotter SEDs at higher $z$’s.
@@
In the left panel of Fig. 5 we plot the peak dust temperatures (derived from best-fit SEDs given in Sect. 7 using Wien’s law) for the star-forming sample of galaxies as a function of the IR luminosity, color-coded with redshift. We can see that, as in the case of Drew & Casey (2022), $T_{\rm d}$ clearly correlates with $L_{\rm IR}$ (even when confining our sample to sources at $z<2$). However, at any given redshift, we do not find any evolution of the dust temperature with $L_{\rm IR}$. If anything, similarly to Schreiber et al. (2018), we can see a minor decline of $T_{\rm d}$ at highest IR luminosities at $z\lesssim 2$. As is discussed in Liang et al. (2019), the differences in dust temperatures are caused mainly by the variations in the specific star formation rate (sSFR; see the right panel of Fig. 7). In this scenario, higher sSFR systems have more young ($t\lesssim 10$ Myrs) star clusters, which heat the dust to higher temperatures, producing more emission in mid-IR wavelengths and hence increasing the values of the effective peak temperatures. This is consistent with the slight decrease in $T_{\rm d}$ at the highest stellar masses and the increase with redshift seen in the left panel of Fig. 5. Since, at any given redshift, the dust temperature variations are small, we plot the mean values of $T_{\rm d}$ as a function of $z$ (black dots) in the right panel of Fig. 5. We find that the data can be well fit with the quadratic equation, where
$T^{\rm MS}_{\rm d}/{\rm K}=(1.13\pm 0.17)\times z^{2}+(0.88\pm 0.78)\times z\\ +(21.74\pm 0.70).$ (10)
@@
For comparison, the results of Schreiber et al. (2018), Viero et al. (2022) and Ismail et al. (2023) are shown as dotted red, dash-dotted green, and dashed blue lines, respectively. It can be seen that our results, as well as the relation of Schreiber et al. (2018), give slightly higher temperatures at low redshifts than Ismail et al. (2023). This is most likely caused by the fact that both in this work and in the work of Schreiber et al. (2018), the emissivity index, $\beta$, has been fixed at constant value (1.96 and 1.5, respectively). As is shown in Ismail et al. (2023), $\beta$ correlates with dust temperature, where warmer SEDs are characterized by lower emissivity values. At low redshifts, where $T_{\rm d}$ is $\sim 20$ K, $\beta$ is expected to vary between 2.5 and 3.5. Fixing $\beta$ at lower values will naturally increase the best-fit dust temperatures. The effect is stronger for lower assumed values of emissivity, which can be seen in the figure. Comparing to Viero et al. (2022), we can see that our results also point toward slightly hotter dust in the early Universe. As is explained in Viero et al. (2022), however, assuming hotter dust at higher redshifts is necessary in order to explain a number of observations, like the lack of dusty objects in the early Universe (Sommovigo et al. 2020), inconsistencies in the high-$z$ IRX-$\beta$ relation (Bouwens et al. 2016), or unusually high dust masses (Bakx et al. 2020).
5.3 Comparison to the literature
Figure 6: Best-fit star-forming MS shown at each redshift bin studied in this work as the solid black line. For comparison, we show the relations of Speagle et al. (2014), Tomczak et al. (2016) and Popesso et al. (2023), plotted with dashed blue, dash-dotted green, and dotted red lines, respectively. While all the functions tend to roughly agree with each other, some inconsistencies can clearly be seen, discussion of which is presented in Sect. 5.3.
@@ 5.3 Comparison to the literature
In Fig. 6, we show the comparison of the SFR-$M_{\ast}$ relation for SFGs (star-forming MS) found in this work (solid black lines) with those found by Speagle et al. (2014), Tomczak et al. (2016) and Popesso et al. (2023), plotted with dashed blue, dash-dotted green, and dotted red lines, respectively. The lines represent the best-fit functions at the central value of each redshift bin studied in this work. The first feature that can be seen is the lack of the turnover for the SFR-$M_{\ast}$ relation of Speagle et al. (2014). As has been pointed out by Popesso et al. (2023), this was likely caused by the fact that the sample studied in Speagle et al. (2014) was mostly confined to sources with stellar masses below $10^{11}\,{\rm M_{\odot}}$.
@@
At low redshifts, for stellar masses ${\rm log}(M_{\ast}/{\rm M_{\odot}})\lesssim 10.5$, our values of the SFR are lower than those of Tomczak et al. (2016) and Popesso et al. (2023). This is most likely caused by fitting our IR data with colder SEDs, where dust temperatures of $\sim 20$ K were assumed (see Fig. 5). At $z>2$ and ${\rm log}(M_{\ast}/{\rm M_{\odot}})\gtrsim 11$, our values of SFR tend to be slightly larger than those of Tomczak et al. (2016) and Popesso et al. (2023). As is explained in Sect. 3.1, all the starburst galaxies that were detected with ALMA in COSMOS and UDS fields were excluded from the sample of this work. However, because the relatively long wavelength of ALMA observations tends to miss hot dust sources (e.g., Chen et al. 2022), we likely failed to identify a fraction of a high-mass starburst population that, due to its high SFRs, would boost the average stacked values. In addition, the determination of the high-mass end of the MS at high redshifts is based on a handful of sources. Looking at Fig. 3 of Popesso et al. (2023), it can be seen that at the highest redshift both the bending and the normalization of the MS are underpredicted by the adopted functional forms.
@@
Figure 7: Logarithm of the sSFR, derived from the best-fit functional forms of the star-forming MS found in this work, plotted as a function of redshift and color-coded with stellar mass. The shape of the relations is consistent with the redshift evolution of the molecular gas fraction in SFGs, as is discussed in Sect. 5.3. For comparison, we show the findings of Popesso et al. (2023) at two stellar mass regimes. The inset plot shown the relation between the sSFR and peak dust temperature. The similar shapes of the sSFR evolution with $z$ and $T_{\rm d}$ is consistent with the scenario, presented in Liang et al. (2019), in which different dust temperature are mainly driven by the variations in the sSFR (see Sect. 5.2).
@@
In Fig. 7, we show the corresponding evolution of the sSFR with redshift, color-coded with the stellar mass. A clear turnover can be seen at $z\sim 2$, which is a direct consequence of the bending of the normalization of the MS with redshift (Eq. 9). This, on the other hand, as is explained in Sect. 5.1, is likely caused by the increase in the molecular gas fraction, which was found by Wang et al. (2022) to be exponentially increasing with redshift. The sSFR-$z$ relation exhibits lower normalization at higher stellar masses, since at masses $\gtrsim M_{0}$ the MS flattens, decreasing the resulting sSFR. In the inset plot, the evolution of the sSFR with the peak dust temperature can be seen. Similar shapes of the sSFR relation with both redshift and dust temperature are consistent with the suggestion (Liang et al. 2019) that the increase in the dust temperature with redshift, found in Sect. 5.2, is indeed caused by the corresponding increase in the sSFRs.
@@
For comparison, the relation of Popesso et al. (2023) for ${\rm log}(M_{\ast}/{\rm M_{\odot}})\simeq 10.0$ and 11.5 are shown as dotted and dashed black lines, respectively. The inconsistencies at low and high redshifts come from different normalizations of the corresponding MSs, as was explained above. In order to investigate those further, in the following section we explore how the evolution of the MS found in this work and the works of Speagle et al. (2014), Tomczak et al. (2016) and Popesso et al. (2023) affects the “growth” of the GSMF.
@@ 5.4 Growth of the galaxy stellar mass function
Figure 8: Comparison between the observed and simulated star-forming GSMFs. The simulated GSMFs are constructed by “growing” the observed function of McLeod et al. (2021), between the redshifts listed in the upper right corner of each panel, according to Eq. 11. The mass evolution resulting from assumed star-forming MSs of Speagle et al. (2014), Tomczak et al. (2016) and Popesso et al. (2023), are plotted with dashed blue, dash-dotted green, and dotted red lines, respectively. The simulated function constructed using the results of this work are presented in solid black, while the thick solid gray line represents the observed GSMF at the final redshift of each simulation. The inconsistencies at low- and high-mass ends of the simulated GSMFs come from differences in the corresponding shapes of the MSs involved (Fig. 6). The significant shift in densities between the observed and simulated GSMFs is caused by the fraction of SFGs, simulated between the redshifts in question, entering a quiescent mode, as is discussed in Sect. 5.4.
@@
Figure 9: Comparison between the observed and simulated GSMFs, where both star-forming and quiescent galaxies are considered. To extend the redshift range, the simulated GSMFs are constructed by “growing” the observed function of Davidzon et al. (2017), between the redshifts listed in the upper right corner of each panel, according to Eq. 11. The mass evolution results from assuming the SFR-$M_{\ast}$ relation characteristic of galaxy samples including both the star-forming and quiescent sources, where we compare our results, shown in solid black, with those of Tomczak et al. (2016), depicted with a dash-dotted green line. The thick solid gray line represents the observed GSMF at the final redshift of each simulation. It can be seen that the significant gap in densities, apparent in Fig. 8, is no longer present, indicating that, indeed, the inconsistencies were caused by the fraction of sources evolving into quiescent mode, as is discussed in Sect. 5.4. The low-mass slope of simulated GSMFs at $z\gtrsim 2$ is inconsistent with the observed function, indicating that the corresponding slope of the SFR-$M_{\ast}$ relations most likely evolves toward steeper values toward earlier epochs.
@@
The star-forming MS effectively quantifies the stellar mass growth of a galaxy with time. We can therefore compare the evolution of the GSMF as implied by any given MS with the observed one. The procedure is described in detail in Leja et al. (2015) and can be summarized as follows. At each redshift, the stellar mass for a given GSMF is being grown assuming the SFR from the MS, causing the GSMF to shift toward higher masses. Additionally, mainly due to winds and outflows, a fraction of the mass is ejected from a stellar population, a careful treatment of which has been presented in Moster et al. (2013). The stellar mass growth with time can therefore be expressed as
@@
We start the procedure with the observed star-forming stellar mass function of McLeod et al. (2021). The stellar mass at any given redshift is evolved according to Eq. 11 assuming different star-forming MSs. In Fig. 8 we show the results of this exercise, with the observed star-forming GSMF at the final redshift shown in gray. The stellar mass function, produced by evolving the observed GSMF between redshifts listed in the upper right corner of each panel, for the MS of Speagle et al. (2014), Tomczak et al. (2016) and Popesso et al. (2023) are plotted with dashed blue, dash-dotted green, and dotted red lines, respectively. The results from this work are shown with a solid black line. It can be seen that the evolved GSMFs differ at both low and high-mass ends, which is due to the differences in the shapes of the corresponding MSs. At low redshifts and stellar masses the MS found in this work predicts lower values of the SFR, which causes the GSMF to grow more slowly in mass and hence the shift toward higher masses is in this case smaller than those produced by Speagle et al. (2014), Tomczak et al. (2016) and Popesso et al. (2023). Similarly, at high redshift and stellar masses, the stellar mass shifts are slightly larger (except for the case of Speagle et al. 2014, where no bending was assumed). The most striking feature of Fig. 8 is the inconsistency of all the evolved functions with the observed GSMF. There are two main reasons for that. Firstly, the mass evolution of the GSMF is going to be affected by galaxy mergers, where mergers are expected to decrease the number density and contribute to the mass growth of galaxies. As is shown by Leja et al. (2015), however, including mergers will change the evolved GSMF at low masses by only $0.1-0.2$ dex between redshifts 2.25 and 0.5. The second reason is the fact that, as GSMF evolves in time, a fraction of SFGs will transform into passive sources, effectively shifting the evolved stellar mass function down toward lower densities.
@@
In order to account for this effect, we additionally evolve the GSMF, where both star-forming and passive galaxies have been included. In order to extend this exercise to higher redshifts, we adopt the stellar mass function of Davidzon et al. (2017), where $z\simeq 6$ has been reached. The SFR-$M_{\ast}$ relation used in this case includes all galaxies (star-forming and quiescent) and is shown in the left panel of Fig. 3. The resulting evolved GSMF at different redshifts is compared to the observed one in Fig. 9 (thin black and thick gray lines, respectively). Since Speagle et al. (2014) and Popesso et al. (2023) did not publish the SFR-$M_{\ast}$ relations for the galaxy samples that include the quiescent sources, we compare our results with Tomczak et al. (2016) only (dashed green line). It is clear from Fig. 9 that the obvious shift in galaxies number density between the observed and simulated functions, apparent in Fig. 8, is no longer present. This confirms that the inconsistency in the star-forming GSMFs is caused by the decrease in density due to the fraction of SFGs transforming into passive sources. For low masses at $z\gtrsim 2$ the low-mass end of the simulated stellar mass functions tends to be steeper than the observed ones. This is caused by the low-mass slope of the observed GSMF undergoing the evolution in redshift toward higher values. In order for our SFR-$M_{\ast}$ relation to produce results consistent with this, the low-mass slope, $\gamma$, from Eq. 7 would also need to evolve. However, as is explained in Sect. 4.2, since our data does not probe low stellar masses at high redshifts, $\gamma$ was assumed to be constant. The fact that the constant $\gamma$ gives steeper low-mass slopes of the GSMF at high $z$’s indicates that the power law slope of the SFR-$M_{\ast}$ relation for all galaxies (star-forming and quiescent) should increase in redshift. Similarly, since the low-mass slope of the star-forming GSMF, found in both Davidzon et al. (2017) and McLeod et al. (2021), increases at $z\gtrsim 2$, the slope of the star-forming MS should also evolve. At $z\gtrsim 3$ the high-mass end of the simulated GSMF overpredicts the number of galaxies. While the errors on the observed high-redshift stellar mass function at these masses are significantly large ($\sim 1$ dex), it is also possible that the high-mass end of the SFR-$M_{\ast}$ relations in this work, and hence the growth of the stellar masses, is contaminated by the presence of starburst galaxies.